OpenStudy (anonymous):

int_{}^{} 1/(x^2+2) dx how to solve it???

OpenStudy (amistre64):

go tan-1 ish on it i believe

OpenStudy (amistre64):

\[\int \frac{1}{x^2+a^2}dx=\frac{1}{a}tan^{-1}(\frac{x}{a})+C\]

OpenStudy (anonymous):

got other ways to solve it??

OpenStudy (amistre64):

nope

OpenStudy (anonymous):

thanks a lot~=)

OpenStudy (amistre64):

yw